# Grammatical Error Correction for Low-Resource Languages: The Case of Zarma

### v1

@@ Abstract
Grammatical error correction (GEC) is important for improving written materials for low-resource languages like Zarma—spoken by over 5 million people in West Africa. Yet it remains a challenging problem. This study compares rule-based methods, machine translation (MT) models, and large language models (LLMs) for GEC in Zarma. We evaluate each approach’s effectiveness on our manually-built dataset of over 250,000 examples using synthetic and human-annotated data. Our experiments show that the MT-based approach using the M2M100 model outperforms others, achieving a detection rate of 95.82% and a suggestion accuracy of 78.90% in automatic evaluations, and scoring 3.0 out of 5.0 in logical/grammar error correction during MEs by native speakers. The rule-based method achieved perfect detection (100%) and high suggestion accuracy (96.27%) for spelling corrections but struggled with context-level errors. LLMs like MT5-small showed moderate performance with a detection rate of 90.62% and a suggestion accuracy of 57.15%. Our work highlights the potential of MT models to enhance GEC in low-resource languages, paving the way for more inclusive NLP tools.
@@
Mamadou K. KEITA1, Christopher Homan3, Sofianou Abdoulaye Hamani1, Adwoa Bremang2, Marcos Zampieri4, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1NILab, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
@@ 1 Introduction
Grammatical error correction (GEC) is an essential task in natural language processing (NLP) that aims to improve the quality and readability of texts by correcting grammatical errors. GEC are important for enhancing written materials, significantly impacting educational outcomes, professional opportunities, and access to information. This is particularly relevant in under-resourced settings where limited access to academic resources and formal training can amplify language proficiency disparities. Furthermore, communities in these settings often rely on local languages for communication and knowledge transmission. While GEC tools are well-developed for high-resource languages—especially English—where extensive annotated datasets and standardized writing systems are available (Napoles et al., 2019), GEC presents significant challenges for low-resource languages.
Zarma, a Nilo-Saharan language spoken by over 5 million people across Niger and neighboring countries (Lewis et al., 2016), exemplifies the difficulties faced by many low-resource languages. The lack of standardized orthography and the limited availability of annotated data make it challenging to develop practical GEC tools. These challenges are not unique to Zarma but are shared by a broad class of low-resource languages, including many African and indigenous languages worldwide. Addressing these challenges requires innovative approaches that can work with minimal data and accommodate non-standardized texts.
The emergence of large language models (LLMs) gave notice to a new era in NLP, empowering machine learning models to learn and generate human-like text across many languages (Devlin et al., 2018; Brown et al., 2020). LLMs exhibit remarkable zero-shot and few-shot learning capabilities (Wan et al., 2023), which are potentially beneficial for low-resource languages. These capabilities allow LLMs to perform effectively with minimal data, making them valuable for languages with limited annotated datasets. However, the use of LLMs is also challenging in low-resource settings. They are primarily trained on data from high-resource languages, which may limit their performance when applied to low-resource languages. The lack of representative training data can lead to errors and biases, reducing their effectiveness. Additionally, the significant computational resources required to fine-tune and deploy LLMs can be a barrier in resource-constrained environments.
This study investigates the potential of LLMs and traditional models to improve GEC for Zarma by comparing conventional rule-based methods, machine translation (MT)-based models, and the novel application of LLMs for Zarma GEC. To explore the performance of these models on other low-resource languages, we replicate our Zarma GEC experiments with Bambara, a West African language spoken in Mali. Through a case study, we highlight the strengths and weaknesses of each method, aiming to bridge the linguistic gap in NLP for low-resource languages like Zarma and Bambara.
Our research addresses the following questions:
@@
Do state-of-the-art LLM models outperform conventional rule-based and MT-based models on GEC for Zarma texts?
@@
What are the specific strengths and limitations of each GEC approach in low-resource settings, considering the lack of standardized orthography and limited annotated data?
@@
  • A comprehensive evaluation of three distinct GEC approaches for Zarma (rule-based, MT-based, and LLM-based). Our findings show that the MT-based approach delivered the highest accuracy with a detection rate of 95.82%, a suggestion accuracy of 78.90%, and acceptable performance in zero-shot scenarios.
  • Reproduction of the experiments with Bambara—a manding family language—to confirm replicability and broaden the study’s scope beyond Zarma.
  • Development and planned public release (upon acceptance) of models fine-tuned for GEC of the tested languages.
@@ 2 Related Work
Researchers have explored the use of LLMs for language-specific tasks, including GEC, demonstrating their adaptability beyond high-resource languages. For instance, a study by Palma Gomez et al. (2023) showed that the MT5 model (Xue et al., 2020), a multilingual transformer model pre-trained on a massive dataset of text and code, could be effectively fine-tuned for GEC in Ukrainian. Song et al. (2023) present an innovative application of LLMs—specifically GPT-4 (OpenAI, 2023)—for generating explanations for grammatical errors. Another promising approach to GEC leverages pre-trained multilingual MT models. The study by (Luhtaru et al., 2024) introduced the "No Error Left Behind" approach, which uses models like MT5 and NLLB (Team et al., 2022). The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence. However, the research also identified a fundamental challenge: the trade-off between precision and recall when training with synthetic data. This finding shows the need for further research to optimize these models for low-resource GEC tasks, potentially through improved data augmentation techniques or developing specialized architectures for error correction.
@@ 3.1 GEC with LLMs
LLMs have significantly advanced NLP, exhibiting capabilities in multitasking, few-shot and multilingual learning. These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020). This section outlines our methodology for leveraging LLMs to develop a GEC tool for Zarma. The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
@@ 3.1.1 Implementation
We employed two distinct approaches for LLM-based GEC:
@@
This method involves embedding training data within a contextual sentence structure, enhancing the model’s reasoning abilities, and facilitating effective learning from the examples. We use the following instruction prompt:
@@
This format, particularly the error explanation component, is important for aiding the model’s comprehension of Zarma’s grammatical contexts and patterns, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
@@
This approach involves fine-tuning the model directly on parallel data of incorrect and correct Zarma sentences without explicit prompts. This leverages the model’s ability to learn the implicit mapping between incorrect and correct forms.
@@ 3.2 GEC with MT Models
Exploring MT models for GEC represents a promising avenue for addressing GEC challenges in low-resource languages like Zarma. MT models, particularly those pre-trained on multilingual datasets, offer strong capabilities for processing text across diverse linguistic frameworks.
@@ 3.2.1 Implementation
We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC. To adapt M2M100 for Zarma GEC, we fine-tuned it on a corrupted corpus designed to reflect common errors in Zarma text. This corpus was generated by applying a custom noise script, described in Section 4. The script introduces various errors, ensuring the training data effectively represents realistic challenges in real-world Zarma text. This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, enabling the model to learn the mapping from incorrect to correct Zarma. Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
@@ 3.3 Rule-based GEC for Zarma
As a baseline, we designed a rule-based GEC process, based on Levenshtein distance and the Bloom filter (Bloom, 1970), for Zarma (Figure 1). Additionally, we implemented a tool and API in Python. The tool—the first of its kind for Zarma, to our knowledge—is designed to cater to a wide range of users. It offers command-line and graphical user interfaces (GUI) and is much less computationally intensive than the LLM-based approach. Moreover, our results show that it provides a competitive spell correction performance compared to the LLMs- and MT-based approaches. To ensure the tool’s accessibility, we plan a public release upon acceptance.
@@
At the core of the system is the trie-based dictionary lookup. If the Bloom Filter decision is positive ($B(T)=\text{True}$), the system proceeds to a Grammatical Check to verify if the word is grammatically correct. If the word is found to be grammatically incorrect, or if $B(T)=\text{False}$ (indicating the word is not in the dictionary), the system performs a Dictionary Lookup to confirm or correct the word.
@@
4 Data Curation Insights
In this section, we detail our approach to gathering and preparing data for training our GEC models for Zarma. We constructed two distinct datasets—a synthetic dataset and a human-annotated “gold” dataset—which we then combined to create a comprehensive training set. The final dataset comprises over 250,000 examples, split into 80% for training, 10% for validation, and 10% for testing.
@@ 4.0.1 Synthetic Data
To address the scarcity of annotated data in Zarma, we generated a synthetic dataset by applying a custom corruption script to the Feriji Dataset (Keita et al., 2024). This script introduces realistic typographical and grammatical errors into Zarma sentences, mimicking plausible human mistakes. The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
Specifically, we applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
@@
The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$. These operations are designed to introduce errors without inadvertently creating another grammatically valid Zarma sentence.
@@ 4.1 Human-Annotated Data
The second dataset, referred to as the ‘gold data,’ was curated by human annotators proficient in Zarma. They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration. This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios. An example from the gold data is shown in Table 2
@@
Table 2: Snapshot of the Gold Data
@@ 5 Experiment
We conducted a series of experiments to evaluate the effectiveness of different GEC methods for Zarma. We selected three models for training based on their demonstrated proficiency in multilingual tasks and aptitude for few-shot learning: Gemma (Team et al., 2024), MT5-small, and M2M100. The training was conducted on Google Cloud, using a Deep Learning Virtual Machine equipped with NVIDIA GPUs T4 x2. Due to computational resource constraints, we applied QLoRA quantization (Dettmers et al., 2023) to the Gemma model to reduce memory requirements. For MT5-small and M2M100, we used smaller variants of the models to ensure feasible training times. For the LLM-based methods (Gemma and MT5-small), we structured the combined dataset as described in Section 4, incorporating both incorrect sentences and their corrections along with error explanations. For the MT-based method (M2M100), we adopted a MT task-specific format, using aligned pairs of incorrect and correct Zarma sentences without explicit error explanations. The detailed training settings for each model are presented in Table 7.
@@ 6 Results & Comparative Analysis
To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted a comprehensive evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers. We included the ME because AE tools are primarily developed and tested on high-resource languages. Therefore, their scores may not reliably reflect performance in low-resource settings like Zarma. The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
@@ 6.1.1 Common Machine Learning Metrics
The GLEU and M2 scores provide an overall assessment of the models’ performance in correcting grammatical errors by comparing the corrected sentences to reference sentences. The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models. The LLM-based models, Gemma 2b and MT5-small, obtained moderate scores, reflecting their ability to correct some errors but not as effectively as the M2M100 model.
@@ 6.1.2 Spelling Correction Metrics
As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%. This impressive performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope. The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%. This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting. The LLM-based models, Gemma 2b and MT5-small, showed lower detection rates and suggestion accuracies. Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%. These results indicate that while LLMs have some capacity to correct spelling errors, they are less effective than the MT-based approach and the rule-based method in this specific task.
@@ 6.3 Zero-Shot scores
To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset containing sentences with words and grammatical structures not present in the training data. This evaluation is crucial for understanding how well the models can handle new or unseen data, which is a common scenario in low-resource language settings. The results of the zero-shot evaluation are presented in Table 5.
@@
The M2M100 model again demonstrated superior performance among the tested models in the zero-shot setting. It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability. Its detection rate was 92.27% with a suggestion accuracy of 78.34%, showing that it can effectively identify and correct errors in unseen data. Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
@@
Our results strongly suggest that MT models trained on extensive multilingual datasets hold significant potential for improving GEC in low-resource languages. The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks. To validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family. The results of this experiment are detailed in Section A.
7 Strengths and Limitations of Each Method
7.1 Rule-Based Method
Rule-based approaches demonstrated high effectiveness in addressing predictable and previously encountered error patterns. In controlled tests, these methods achieved perfect detection and high suggestion accuracy. Their strength lies in the reliance on a comprehensive set of predefined rules and a detailed lexicon. However, this reliance also presents significant limitations—inflexibility and lack of adaptability. Rule-based methods struggle to handle new or unexpected errors, especially in dynamic language use where novel words or grammatical constructions are common. This limitation is particularly pronounced in zero-shot scenarios, where the system encounters words or structures not included in its rules or dictionary. Moreover, the dependency on extensive and carefully curated linguistic resources restricts the scalability of rule-based methods, especially for low-resource languages like Zarma, where such resources are often limited or incomplete (Scannell, 2007).
7.2 LLMs Method
The LLMs in our experiments—Gemma 2b and MT5-small—demonstrated moderate performance in both controlled and zero-shot scenarios. A key strength of LLMs is their capacity to understand context and generate corrections based on broader textual cues, rather than relying only on direct matches to known errors. However, their performance is significantly dependent on the diversity and quality of the training data. A major limitation is that most pre-existing LLMs are primarily trained on data from high-resource languages—mainly Western languages. Consequently, their applicability to African languages like Zarma is hindered by a lack of representative training examples (Bender et al., 2021). This results in lower suggestion accuracy and difficulties in effectively handling the linguistic complexities of these languages. Furthermore, training and fine-tuning LLMs require substantial computational resources, creating a significant barrier in resource-constrained environments.
7.3 MT Method
The MT approach—using the M2M100 model—demonstrated exceptional performance, particularly in zero-shot scenarios, surpassing both rule-based methods and LLMs. The strength of this approach lies in the ability of MT models to leverage multilingual translation mechanisms, effectively adapting to the nuances of diverse languages through exposure to vast and varied training datasets. This characteristic makes MT models particularly suitable for GEC in low-resource languages, as they can infer corrections from patterns learned across multiple languages. This aligns with research highlighting the effectiveness of MT models for cross-lingual transfer learning in various NLP tasks (Conneau et al., 2018). However, a significant challenge in using MT models for GEC in low-resource languages is the frequent scarcity of quality parallel corpora for training. Without sufficient data, the models may generate less accurate or contextually inappropriate corrections (Tiedemann, 2020). Additionally, despite their strengths, MT models require fine-tuning and continuous updating to maintain their accuracy and relevance, as language use evolves.

### v2

@@ Abstract
Grammatical error correction (GEC) aims to improve quality and readability of texts through accurate correction of linguistic mistakes. Previous work has focused on high-resource languages, while low-resource languages lack robust tools. However, low-resource languages often face problems such as: non-standard orthography, limited annotated corpora, and diverse dialects, which slows down the development of GEC tools. We present a study on GEC for Zarma, spoken by over five million in West Africa. We compare three approaches: rule-based methods, machine translation (MT) models, and large language models (LLMs). We evaluated them using a dataset of more than 250,000 examples, including synthetic and human-annotated data. Our results showed that the MT-based approach using M2M100 outperforms others, with a detection rate of 95. 82% and a suggestion accuracy of 78. 90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME) from native speakers for grammar and logical corrections. The rule-based method was effective for spelling errors but failed on complex context-level errors. LLMs—MT5-small—showed moderate performance. Our work supports use of MT models to enhance GEC in low-resource settings, and we validated these results with Bambara, another West African language.
@@
Mamadou K. KEITA1, Christopher Homan3, Marcos Zampieri4, Adwoa Bremang2, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1PAAIR Institute, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
@@ 1 Introduction
Grammatical error correction (GEC) is important to improving text quality in different scenarios. High-resource languages have well-developed GEC tools (Napoles et al., 2019), but low-resource languages have few resources. Zarma, a Nilo-Saharan language spoken widely in West Africa (Lewis et al., 2016), faces these challenges.
Zarma texts lack standardized orthography and large annotated datasets. Developers need approaches that handle minimal resources and non-standard writing. Large language models (LLMs) hold promise—as they can handle various languages and tasks (Brown et al., 2020)—but performance on low-resource languages can be limited. They rely heavily on data, which is scarce for Zarma.
We investigate three GEC approaches for Zarma: a rule-based method, a machine translation (MT)-based model, and LLMs. We also test with Bambara, another West African language, to confirm generalization. This paper addresses two key questions:
@@
Do LLMs outperform traditional rule-based and MT models in Zarma GEC?
@@
What are the strengths and limitations of each approach, given the non-standard orthography and limited data?
RQ3:
Does the best-performing approach replicate for another low-resource language—Bambara?
@@
  • A comparison of three GEC methods for Zarma. The MT-based approach—M2M100—provided the best results, with a detection rate of 95. 82% and a suggestion accuracy of 78. 90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME).
  • A dataset of more than 250,000 synthetic and human-annotated Zarma examples.
  • Experiments with Bambara to confirm replicability.
  • Planned release of trained models to promote future research.
@@ 2 Related Work
Researchers have explored the use of LLMs for language-specific tasks, including GEC, demonstrating their adaptability beyond high-resource languages. For instance, a study by Palma Gomez et al. (2023) showed that the MT5 model (Xue et al., 2020), a multilingual transformer model pre-trained on a massive dataset of text and code, could be effectively fine-tuned for GEC in Ukrainian. Song et al. (2023) present an innovative application of LLMs—specifically GPT-4 (OpenAI, 2023)—for generating explanations for grammatical errors. Another promising approach to GEC leverages pre-trained multilingual MT models. The study by (Luhtaru et al., 2024) introduced the "No Error Left Behind" approach, which uses models like MT5 and NLLB (Team et al., 2022). The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence. The complexity of balancing precision and recall on synthetic data remains an open problem.
@@ 3.1 GEC with LLMs
LLMs have significantly advanced NLP, showing capabilities in multitasking, single-shot, and multilingual learning. These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020). This section outlines our methodology for using LLMs to develop a GEC tool for Zarma. The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
@@ 3.1.1 Implementation
We used two distinct approaches for LLM-based GEC:
@@
This method involves embedding training data within a contextual sentence structure. We embed examples in a prompt format:
@@
This format helps the model reason about corrections and reflect on why errors occur, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
@@
We also fine-tuned LLMs on parallel incorrect and correct sentences. This leverages the model’s ability to align corrupted text with corrected forms.
@@ 3.2 GEC with MT Models
Using MT models for GEC represents a promising avenue for addressing GEC challenges in low-resource languages such as Zarma. MT models—particularly those pre-trained on multilingual datasets—offer strong capabilities for processing text across diverse linguistic frameworks.
@@ 3.2.1 Implementation
We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC. We adapt M2M100 for GEC by training it to translate corrupted Zarma text into corrected versions. This approach uses a synthetic corpus generated with a noise script (Section 4). The script introduces various errors, ensuring that the training data effectively represent realistic challenges in real-world Zarma text. This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, allowing the model to learn the mapping from incorrect to correct Zarma. Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
@@ 3.3 Rule-based GEC for Zarma
As a baseline, we built a rule-based system using Levenshtein distance and a Bloom filter (Bloom, 1970), for Zarma (Figure 1). Additionally, we implemented a tool and API in Python. The tool—the first of its kind for Zarma, to our knowledge—is designed for a wide range of users. It offers command-line and graphical user interfaces (GUI) and is much less computationally intensive than the LLM-based approach. Moreover, our results show that it provides a competitive spell correction performance compared to the LLM- and MT-based approaches. To ensure the accessibility of the tool, we plan a public release upon acceptance.
@@
At the core of the system is the trie-based dictionary lookup. We built this dictionary using the Feriji dataset dictionary (Keita et al., 2024). If the Bloom Filter decision is positive ($B(T)=\text{True}$), the system proceeds to a Grammatical Check to verify if the word is grammatically correct. If the word is found to be grammatically incorrect, or if $B(T)=\text{False}$ (indicating the word is not in the dictionary), the system performs a Dictionary Lookup to confirm or correct the word.
@@
4 Data
We created a combined dataset of over 250,000 examples by merging two distinct datasets—a synthetic dataset and a human-annotated ”gold” dataset. We split it into 80% training, 10% validation, and 10% testing.
@@ 4.0.1 Synthetic Data
To address the scarcity of annotated data in Zarma, We applied a noise script to the Feriji dataset (Keita et al., 2024). This script introduces realistic typographical and grammatical errors into Zarma sentences, to mimick plausible human mistakes. The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
We applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
@@
The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$.
@@ 4.1 Human-Annotated Data
The second dataset was curated by human annotators proficient in Zarma. They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration. This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios. An example from the human-annotated dataset is shown in Table 2
@@
Table 2: Snapshot of the human-annotated dataset
@@ 5 Experiment
We conducted a series of experiments to evaluate the effectiveness of different GEC methods for Zarma. We selected three models for training based on their demonstrated proficiency in multilingual tasks and aptitude for few-shot learning: Gemma (Team et al., 2024), MT5-small, and M2M100. The training was conducted on Google Cloud, using a Deep Learning Virtual Machine equipped with NVIDIA GPUs T4 x2. Due to computational resource constraints, we applied QLoRA quantization (Dettmers et al., 2023) to the Gemma model to reduce memory requirements. For MT5-small and M2M100, we used smaller variants of the models to ensure feasible training times. For LLM-based methods (Gemma and MT5-small), we structured the combined dataset as described in Section 4, to incorporate both incorrect sentences and their corrections along with error explanations. For the MT-based method (M2M100), we adopted a MT task-specific format, using aligned pairs of incorrect and correct Zarma sentences without explicit error explanations. The detailed training settings for each model are presented in Table 7.
@@ 6 Results & Comparative Analysis
To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted an evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers. We included the ME because AE tools are primarily developed and tested on high-resource languages. Therefore, their scores may not reliably reflect performance in low-resource settings such as Zarma. The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
@@ 6.1.1 Common Machine Learning Metrics
The GLEU and M2 scores provide an overall assessment of the models’ performance in correcting grammatical errors by comparing the corrected sentences to reference sentences. The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models. The LLM-based models—Gemma 2b and MT5-small—obtained moderate scores, which shows their ability to correct some errors but not as effectively as the M2M100 model.
@@ 6.1.2 Spelling Correction Metrics
As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%. This high performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope. The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%. This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting. The LLM-based models—Gemma 2b and MT5-small—showed lower detection rates and suggestion accuracies. Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%. These results indicate that while LLMs have some capacity to correct spelling errors, they are less effective than the MT-based approach and the rule-based method in this specific task.
@@ 6.3 Zero-Shot scores
To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset. For that, We gathered a separate set of Zarma texts from books that had not been used during training. We then applied the same corruption script described in Section 4 and asked the annotators to include some LG errors. containing sentences with words and grammatical structures not present in the training data. This evaluation aimed to understand how well the models can handle new or unseen data, which is a common scenario in low-resource language settings. The results of the zero-shot evaluation are presented in Table 5.
@@
The M2M100 model again demonstrated higher performance among the tested models in the zero-shot setting. It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability. Its detection rate was 92.27% with a suggestion accuracy of 78.34%, showing that it can effectively identify and correct errors in unseen data. Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
@@
Our results suggest that MT models trained on multilingual datasets hold significant potential for improving GEC in low-resource languages. The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks. To validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family. The results of this experiment are detailed in Section A.